# Linking emitted drops to collective bursting bubbles across a wide range of bubble size distributions

### v1

@@ Abstract
Bubbles entrained by breaking waves rise to the ocean surface, where they cluster before bursting and release droplets into the atmosphere. The ejected drops and dry aerosol particles, left behind after the liquid drop evaporates, affect the radiative balance of the atmosphere and can act as cloud condensation nuclei. The remaining uncertainties surrounding the sea spray emissions function motivate controlled laboratory experiments that directly measure and link collective bursting bubbles and the associated drops and sea salt aerosols. We perform experiments in artificial seawater for a wide range of bubble size distributions, measuring both bulk and surface bubble distributions (measured radii from 30 µm to 5 mm), together with the associated drop size distribution (salt aerosols and drops of measured radii from 50 nm to 500 µm) to quantify the link between emitted drops and bursting surface bubbles. We evaluate how well the individual bubble bursting scaling laws describe our data across all scales and demonstrate that the measured drop production by collective bubble bursting can be represented by a single framework integrating individual bubble bursting scaling laws over the various bubble sizes present in our experiments. We show that film drop production by bubbles between 100 µm and 1 mm describes the submicron drop production, while jet drop production by bubbles from 30 µm to 2 mm describes the production of drops larger than 1 µm. Our work confirms that sea spray emissions functions based on individual bursting processes are reasonably accurate as long as the surface bursting bubble size distribution is known.
@@ 1 Introduction
When a wave breaks at sea, it entrains air bubbles below the ocean surface. Those bubbles rise back to the surface and reside for some time before bursting and ejecting drops into the atmosphere. The ejected drops can then evaporate, leaving behind ocean spray aerosol particles composed of salts, biological material, and other chemicals found in the ocean. In the atmosphere, these emitted drops and corresponding dry particles affect the radiative balance and can serve as cloud condensation nuclei (Lewis and Schwartz,, 2004; de Leeuw et al.,, 2011; Veron,, 2015; Cochran et al.,, 2017).
@@
One main difficulty of studying spray generation by bubble bursting is the range of scales involved in the related processes, from the single breaking wave of O(1–10 m) to entrained bubbles of sizes O(0.01–10 mm) and emitted drops of sizes O(0.01–100 µm). To predict ocean spray generation from knowledge of the bubble size distribution, we must understand the physical link between underwater bubbles that rise to the ocean surface, surface bubbles that cluster and burst, and the resulting drops they eject (depicted in figure 1 of Deike, (2022)).
@@
Parameterizations of the sea spray generation function used in large-scale models (as reviewed by Lewis and Schwartz, (2004); de Leeuw et al., (2011); Veron, (2015)), display large uncertainties of at least an order of magnitude across the full range of emitted drop sizes. These uncertainties come from the difficulty of characterizing the drop formation across many processes and scales, as well as from the various empirical ways the parameterizations are developed from interpretation of field measurements or upscaling of laboratory data. In practice, the uncertainties lead to very different sensitivity to physical parameters such as temperature (Forestieri et al.,, 2018) and directly impact the modeled radiative balance and climate sensitivity in climate models (Paulot et al.,, 2020).
@@
Individual bubble bursting has been extensively studied to characterize the production mechanism, size, and number of drops produced by the bursting of a single bubble. Previous laboratory experiments have focused either on film drops (Blanchard and Syzdek,, 1988; Resch and Afeti,, 1991, 1992; Spiel, 1997a, ; Spiel,, 1998; Lhuissier and Villermaux,, 2012; Jiang et al.,, 2022) or jet drops (Spiel, 1997a, ; Spiel,, 1998; Ghabache et al.,, 2014; Ghabache and Séon,, 2016; Walls et al.,, 2015; Brasz et al.,, 2018), and numerical simulations have studied jet drop production (Deike et al.,, 2018; Brasz et al.,, 2018; Berny et al.,, 2020, 2021). From those studies, scaling laws have been developed that relate the bursting bubble size to the number and size of drops emitted through jet drop production (Gañán-Calvo,, 2017, 2023; Blanco–Rodríguez and Gordillo,, 2020), film drop production through a centrifuge mechanism (Lhuissier and Villermaux,, 2012), and film drop production through a proposed flapping mechanism (Jiang et al.,, 2022, 2024). However, the extent to which these scaling laws, developed for idealized individual bubble bursting, apply to more realistic configurations like clusters of bubbles within ocean whitecaps, remains to be experimentally demonstrated.
@@
Combining the above-mentioned individual bubble bursting scaling laws for the various mechanisms with the bubble size distribution under a breaking wave and the statistics of air entrainment at large scale, Deike et al., (2022) developed a mechanistic sea spray emissions function and demonstrated the remarkable coherence with empirically-derived emissions functions. To obtain the ocean spray emissions function, Deike et al., (2022) integrates over a given bubble size distribution (constrained by laboratory and numerical studies by Deane and Stokes, (2002); Deike et al., (2016); Mostert et al., (2022)), assuming individual bubble bursting scaling laws for the size, number, and distribution of emitted drops for each production mechanism. The method uses the integration introduced for film drop production by Lhuissier and Villermaux, (2012) and extended to jet drop production in Berny et al., (2021), and air entrainment volume by breaking waves statistics in the ocean (Deike et al.,, 2016, 2017).
@@
The approach from Deike et al., (2022), as well as past approaches (see the overview in Lewis and Schwartz, (2004)), assumes that the bubble size distribution under a breaking wave is always of the same shape and does not consider collective bubble bursting or surfactant effects. Physico-chemical properties of the ocean water – such as salinity, temperature, surfactant concentration, and biological composition – bring another layer of complexity to sea spray generation by potentially modulating the bubble processes (i.e. lifetime, coalescence, bursting). Modeling of sea spray organic aerosol emissions in Burrows et al., (2014) highlights the numerous chemical and biological factors that are thought to influence spray generation, showing that – despite extensive past experimental efforts – there is still a need for controlled, well-constrained experiments to improve our knowledge of how these realistic factors control the bursting processes while limiting experimental uncertainties.
@@
To investigate the collective bursting relevant to the open ocean context, previous experiments have measured drop production by collections of bubbles generated through different methods in a variety of complex solutions, focusing especially on the measurement of submicron drops/aerosols. The experiments include bursting bubbles generated by both a plunging sheet and nucleation bubbler in natural and artificial seawater (Wang et al.,, 2017), bursting bubble rafts produced in sampled natural seawater through a forced-air venturi channel (Frossard et al.,, 2019), the investigation of the effect of phytoplankton and bacteria on aerosol production for bubbles generated by laboratory breaking waves in natural seawater (Prather et al.,, 2013), as well as several other experiments involving bursting bubble rafts or foams in solutions of different organic content, or different salt concentration (Cipriano and Blanchard,, 1981; Mårtensson et al.,, 2003; Sellegri et al.,, 2006; Tyree et al.,, 2007; Modini et al.,, 2013; Zinke et al.,, 2022; Dubitsky et al.,, 2023). These studies suggest that ocean physico-chemical parameters modulate the bursting process. However, the complexity of the seawater solution and the lack of complete bulk bubble, surface bubble, and drop/dry particle measurements across a wide range of scales make it difficult to compare results from separate studies and to distinguish the specific effects of individual physico-chemical variables among other differences due to varied bubble size distributions or collective bursting.
@@
These uncertainties and their practical consequences on ocean spray emissions functions motivate an intermediate-scale collective bursting laboratory experiment, measuring bubbles in the bulk and at the surface as well as the produced drops, as a way to bridge the idealized and the oceanic spray generation configurations. Using a well-characterized setup with measurements that capture the full size range of bubbles and associated ejected drops, we can maintain precise control over the solution while observing the effect of any changes in the properties on the collective bursting. This idea was introduced through the experiments of Néel and Deike, (2021); Néel et al., (2022), where systematic measurements of millimetric bubbles and supermicron drops were made for solutions of different surfactant concentrations.
@@
In this paper, we perform experiments with measurements of all relevant scales of bulk and surface bubbles (radii 30 µm to 5 mm), liquid drops (radii 10 to 500 µm), and salt aerosols (equivalent drop radii 50 nm to 10 µm) for various clustered bubble configurations, allowing us to directly characterize the link between the drop size distribution and the surface bursting bubble distribution. We have designed a bubbling tank experiment (building off the work from Néel and Deike, (2021)) where bubbles are generated at the bottom of the tank, and statistical measurements are made in the bulk and at the surface. Three measurement techniques are combined to measure drop sizes spanning four orders of magnitude: optical digital inline holography allows the measurement of drops ejected in the air with radii from 10 µm to 500 µm, while smaller droplets (not accessible with direct imaging techniques) are inferred by drying and measuring the size of the resulting salt crystals, spanning equivalent drop radii from 50 nm to 10 µm. We characterize the spray generation by clusters of bubbles in a precisely-controlled solution of artificial seawater, and we vary the bulk bubble size distribution across cases. The bulk bubble size distributions tested include quasi-monodisperse plumes of rising bubbles centered around 3 mm, 2 mm, and 250 µm, as well as broad-banded distributions where bubbles range from 30 µm to 3 mm. The broad-banded size distributions mimic the bubble size distribution entrained by breaking waves. As illustrated in figure 1, these bubble size distributions exhibit very different characteristics at the surface, which will control the resulting drop size distribution. We then explore whether we can describe the associated spray generation by these various bubble distributions within a single physical framework.
@@ 2.1 Experimental setup and conditions
Figure 2: Sketch of the experimental setup, with measurements of bubbles, drops, and dry aerosol particles. Bubbles of nearly identical sizes are generated by compressed air flowing through needles at the bottom of the tank. When two underwater pumps are turned on (as illustrated in the sketch), the injected bubbles are broken up into many smaller bubbles as they rise into the region of turbulence generated by the water jets from the pumps (bubbles not drawn to scale). The bubbles are imaged from the side view to measure the rising bulk bubbles, and from the top view to capture the bubbles at the surface. Large fields of view (black) and small fields of view (orange) are drawn approximately to scale and permit measurement of bubble radii from 30 µm to 5 mm. Liquid drops ejected into the air by bursting bubbles are measured through an inline holographic setup (measuring liquid drop radii from 10µm to 500 µm), while dry aerosol particles are measured by extracting and drying the liquid drops and analyzing them using an Optical Particle Sizer and Scanning Mobility Particle Sizer (OPS, SMPS composed of DMA and CPC - TSI, Inc.) (for dry particle radii from 25 nm to 5 µm). A line of filtered, compressed air is connected to the lid to flush the air above the water surface before the start of each run. Relative humidity and both air and water temperature are monitored throughout the experiments.
@@
The experimental setup shown in figure 2, consists of a polycarbonate tank with inner dimensions 47 x 47 x 61 cm3, similar to the one used by Néel and Deike, (2021), which is filled to a height of 35 cm with a solution of artificial seawater (made of deionized (DI) water and artificial sea salt, ASTM D1141-98, Lake Products Company LLC). The water height is chosen so that the waves/fluctations on the water surface remain small and the surface bubbles can be clearly imaged. Bubbles are continuously generated as compressed air flows through the needles positioned at the bottom of the tank, arranged in a square with eight needles extending horizontally from each side. The air flowrate through the needles is controlled at 100 cm3/min per needle, so that the resulting plume of rising bubbles is fairly monodisperse.
@@
The artificial seawater solution is prepared by mixing DI water with ASTM D1141-98 Artificial Sea Salt, which contains the same proportions of elements found in natural seawater in amounts greater than 0.0004% by weight (Lake Products Company LLC). Solutions discussed in this paper were prepared with a salinity of 36 g/L (35 g/kg) to mimic the composition of ocean water. The solution has liquid viscosity $\mu$ = 1.001 mPa s, liquid density $\rho$ = 1.025 g/m3, and static surface tension $\gamma$ = 52 mN/m. The low static surface tension indicates that the salt mixture likely includes some surfactant. At the beginning and end of each day of experiments, the solution was characterized by measuring the surface tension isotherm with a Langmuir Trough (KSV NIMA, model KN 1003). This measurement was used to confirm that the surface properties of the solution remained consistent over multiple days of experiments.
Two pumps (Rule 20DA 800 GPH Bilge Pump) are mounted 6 cm above the needle tips, pointed towards each other from opposite corners. When the pumps are turned on, they produce water jets that meet in the center of the tank, creating a region of turbulent flow that causes the rising bubbles to fragment into many smaller bubbles, with turbulence similar to the setup described in Ruth et al., (2022). Experiments can be performed with or without forced turbulence, where the turbulence has two effects: agitate the flow and break the bubbles, leading to a broad-banded size distribution. Note that bubble imaging occurs above the region of active breakup.
@@
Case Description Bubble Needle ID Pumps # Needles $R_{b,inj}$ $N_{R_{b}<2mm}$
Production Method (µm) (mm) (cm-3)
Narrow - 20mm Needles 432 off 32 2 0.40
Broad - 20mm Needles 432 on 32 2 45
Broad - half volume Needles 432 on 16 2 25
Narrow - 30mm Needles 965 off 32 3 0.13
Broad - 30mm Needles 965 on 32 3 43
Narrow - 2500µm Porous Media (Bubbler) off 0.25 57
Table 1: Experimental parameters. The bulk bubble size distribution across cases was varied by changing between quiescent and turbulent flow underwater (by turning the underwater pumps off or on, causing the bubbles to break up in the turbulent case), by changing the injection bubble size (using two different needle sizes or an aquarium bubbler), and by halving the volume (using 16 instead of 32 needles). The flowrate of air used to create the bubbles was 1000cm3/min per needle for all needle cases, and 8000cm3/min total for the bubbler case. Case names describe the range of radii over which a significant number of bubbles are concentrated (narrow, broad) and the initial injection bubble size before any breakup in turbulence (20mm, 30mm, 2500µm). The broad half volume case is characterized by bubbles arriving at the surface isolated from each other, bursting individually instead of clustered as the smaller number of needles leads to both reduced volume and increased needle spacing. $R_{b,inj}$ represents the typical injection bubble size by the needles or bubbler. $N_{R_{b}<2mm}$ shows the number of bubbles (per unit volume) smaller than radius 20mm, which reinforces the designation of certain cases as narrow- or broad-banded. Bubble size distributions for each case are shown throughout sections 2.2.4, 4.1, and 4.2. Representative surface bubble images for each case are shown in figure 1.
@@
In order to capture the large range of length scales present in the system – spanning bubble radii of 30 to 5000 µm and drop sizes of 0.05 to 500 µm – we make multiple measurements of bulk bubbles, surface bubbles, drops, and dry particles. Bulk bubbles are imaged in both a large and small field-of-view by cameras positioned above the region of active breakup. Surface bubbles are also captured in two fields-of-view by cameras positioned perpendicular to the water surface, where the bubbles cluster near the center of the tank before bursting and ejecting drops into the air above the water surface, which is enclosed in the tank with an acrylic lid.
Drops are sized using a combination of an inline holographic system positioned just above the water surface Erinin et al., (2019, 2023), an Optical Particle Sizer (OPS: TSI, Inc. OPS 3330), and a Scanning Mobility Particle Sizer (SMPS, consists of: TSI, Inc. Advanced Aerosol Neutralizer, Model 3088; Electrostatic Classifier, Model 3082; Differential Mobility Analyzer (DMA), Model 3081; Condensation Particle Counter (CPC), Model 3752). The OPS and SMPS are connected through the lid and pull air with suspended drops out of the tank, where they are dried and then measured. Relative humidity and air temperature are monitored 25 cm above the calm water surface throughout all experimental runs, along with the temperature of the solution.

### v2

@@ Abstract
Bubbles entrained by breaking waves rise to the ocean surface, where they cluster before bursting and release droplets into the atmosphere. The ejected drops and dry aerosol particles, left behind after the liquid drop evaporates, affect the radiative balance of the atmosphere and can act as cloud condensation nuclei. The remaining uncertainties surrounding the sea spray emissions function motivate controlled laboratory experiments that directly measure and link collective bursting bubbles and the associated drops and sea salt aerosols. We perform experiments in artificial seawater for a wide range of bubble size distributions, measuring both bulk and surface bubble distributions (measured radii from 30 µm to 5 mm), together with the associated drop size distribution (salt aerosols and drops of measured radii from 50 nm to 500 µm) to quantify the link between emitted drops and bursting surface bubbles. We evaluate how well the individual bubble bursting scaling laws describe our data across all scales and demonstrate that the measured drop production by collective bubble bursting can be represented by a single framework integrating individual bursting scaling laws over the various bubble sizes present in our experiments. We show that film drop production by bubbles between 100 µm and 1 mm describes the submicron drop production, while jet drop production by bubbles from 30 µm to 2 mm describes the production of drops larger than 1 µm. Our work confirms that sea spray emissions functions based on individual bursting processes are reasonably accurate as long as the surface bursting bubble size distribution is known.
@@ 1 Introduction
When a wave breaks at sea, it entrains air bubbles below the ocean surface. Those bubbles rise back to the surface and reside for some time before bursting and ejecting drops into the atmosphere. The ejected drops can then evaporate, leaving behind ocean spray aerosol particles composed of salts, biological material, and other chemicals found in the ocean. In the atmosphere, these emitted drops and corresponding dry particles affect the radiative balance and can serve as cloud condensation nuclei (Lewis & Schwartz, 2004; de Leeuw et al., 2011; Veron, 2015; Cochran et al., 2017).
@@
One main difficulty of studying spray generation by bubble bursting is the range of scales involved in the related processes, from the single breaking wave of O(1–10 m) to entrained bubbles of sizes O(0.01–10 mm) and emitted drops of sizes O(0.01–100 µm). To predict ocean spray generation from knowledge of the bubble size distribution, we must understand the physical link between underwater bubbles that rise to the ocean surface, surface bubbles that cluster and burst, and the resulting drops they eject (depicted in figure 1 of Deike (2022)).
@@
Parameterizations of the sea spray generation function used in large-scale models (as reviewed by Lewis & Schwartz (2004); de Leeuw et al. (2011); Veron (2015)), display large uncertainties of at least an order of magnitude across the full range of emitted drop sizes. These uncertainties come from the difficulty of characterizing the drop formation across many processes and scales, as well as from the various empirical ways the parameterizations are developed from interpretation of field measurements or upscaling of laboratory data. In practice, the uncertainties lead to very different sensitivity to physical parameters such as temperature (Forestieri et al., 2018) and directly impact the modeled radiative balance and climate sensitivity in climate models (Paulot et al., 2020).
@@
Individual bubble bursting has been extensively studied to characterize the production mechanism, size, and number of drops produced by the bursting of a single bubble. Previous laboratory experiments have focused either on film drops (Blanchard & Syzdek, 1988; Resch & Afeti, 1991, 1992; Spiel, 1997a, 1998; Lhuissier & Villermaux, 2012; Jiang et al., 2022) or jet drops (Spiel, 1997a, 1998; Ghabache et al., 2014; Ghabache & Séon, 2016; Walls et al., 2015; Brasz et al., 2018), and numerical simulations have studied jet drop production (Deike et al., 2018; Brasz et al., 2018; Berny et al., 2020, 2021). From those studies, scaling laws have been developed that relate the bursting bubble size to the number and size of drops emitted through jet drop production (Gañán-Calvo, 2017, 2023; Blanco–Rodríguez & Gordillo, 2020), film drop production through a centrifuge mechanism (Lhuissier & Villermaux, 2012), and film drop production through a proposed flapping mechanism (Jiang et al., 2022, 2024). However, the extent to which these scaling laws, developed for idealized individual bubble bursting, apply to more realistic configurations like clusters of bubbles within ocean whitecaps, remains to be experimentally demonstrated.
@@
Combining the above-mentioned individual bubble bursting scaling laws for the various mechanisms with the bubble size distribution under a breaking wave and the statistics of air entrainment at large scale, Deike et al. (2022) developed a mechanistic sea spray emissions function and demonstrated the remarkable coherence with empirically-derived emissions functions. To obtain the ocean spray emissions function, Deike et al. (2022) integrates over a given bubble size distribution (constrained by laboratory and numerical studies by Deane & Stokes (2002); Deike et al. (2016); Gao et al. (2021); Mostert et al. (2022)), assuming individual bubble bursting scaling laws for the size, number, and distribution of emitted drops for each production mechanism. The method uses the integration introduced for film drop production by Lhuissier & Villermaux (2012) and extended to jet drop production in Berny et al. (2021), and air entrainment volume by breaking waves statistics in the ocean (Deike et al., 2016, 2017).
@@
The approach from Deike et al. (2022), as well as past approaches (see the overview in Lewis & Schwartz (2004)), assumes that the bubble size distribution under a breaking wave is always of the same shape and does not consider collective bubble bursting or surfactant effects. Physico-chemical properties of the ocean water – such as salinity, temperature, surfactant concentration, and biological composition – bring another layer of complexity to sea spray generation by potentially modulating the bubble processes (i.e. lifetime, coalescence, bursting). Modeling of sea spray organic aerosol emissions in Burrows et al. (2014) highlights the numerous chemical and biological factors that are thought to influence spray generation, showing that – despite extensive past experimental efforts – there is still a need for controlled, well-constrained experiments to improve our knowledge of how these realistic factors control the bursting processes while limiting experimental uncertainties.
@@
To investigate the collective bursting relevant to the open ocean context, previous experiments have measured drop production by collections of bubbles generated through different methods in a variety of complex solutions, focusing especially on the measurement of submicron drops/aerosols. The experiments include bursting bubbles generated by both a plunging sheet and nucleation bubbler in natural and artificial seawater (Wang et al., 2017), bursting bubble rafts produced in sampled natural seawater through a forced-air venturi channel (Frossard et al., 2019), the investigation of the effect of phytoplankton and bacteria on aerosol production for bubbles generated by laboratory breaking waves in natural seawater (Prather et al., 2013), as well as several other experiments involving bursting bubble rafts or foams including those in solutions of different organic content, or different salt concentration (Cipriano & Blanchard, 1981; Mårtensson et al., 2003; Liger-Belair & Jeandet, 2003; Sellegri et al., 2006; Tyree et al., 2007; Modini et al., 2013; Séon & Liger-Belair, 2017; Zinke et al., 2022; Dubitsky et al., 2023). These studies suggest that ocean physico-chemical parameters modulate the bursting process. However, the complexity of the seawater solution and the lack of complete bulk bubble, surface bubble, and drop/dry particle measurements across a wide range of scales make it difficult to compare results from separate studies and to distinguish the specific effects of individual physico-chemical variables among other differences due to varied bubble size distributions or collective bursting.
@@
These uncertainties and their practical consequences on ocean spray emissions functions motivate an intermediate-scale collective bursting laboratory experiment, measuring bubbles in the bulk and at the surface as well as the produced drops, as a way to bridge the idealized and the oceanic spray generation configurations. Using a well-characterized setup with measurements that capture the full size range of bubbles and associated ejected drops, we can maintain precise control over the solution while observing the effect of any changes in the properties on the collective bursting. This idea was introduced through the experiments of Néel & Deike (2021); Néel et al. (2022), where systematic measurements of millimetric bubbles and supermicron drops were made for solutions of different surfactant concentrations.
@@
In this paper, we perform experiments with measurements of all relevant scales of bulk and surface bubbles (radii 30 µm to 5 mm), liquid drops (radii 10 to 500 µm), and salt aerosols (equivalent drop radii 50 nm to 10 µm) for various clustered bubble configurations, allowing us to directly characterize the link between the drop size distribution and the surface bursting bubble distribution. We have designed a bubbling tank experiment (building off the work from Néel & Deike (2021)) where bubbles are generated at the bottom of the tank, and statistical measurements are made in the bulk and at the surface. Three measurement techniques are combined to measure drop sizes spanning four orders of magnitude: optical digital inline holography allows the measurement of drops ejected in the air with radii from 10 µm to 500 µm, while smaller droplets (not accessible with direct imaging techniques) are inferred by drying and measuring the size of the resulting salt crystals, spanning equivalent drop radii from 50 nm to 10 µm. We characterize the spray generation by clusters of bubbles in a precisely-controlled solution of artificial seawater, and we vary the bulk bubble size distribution across cases. The bulk bubble size distributions tested include quasi-monodisperse plumes of rising bubbles centered around 3 mm, 2 mm, and 250 µm, as well as broad-banded distributions where bubbles range from 30 µm to 3 mm. The broad-banded size distributions mimic the bubble size distribution entrained by breaking waves. As illustrated in figure 1, these bubble size distributions exhibit very different characteristics at the surface, which will control the resulting drop size distribution. We then explore whether we can describe the associated spray generation by these various bubble distributions within a single physical framework.
@@ 2.1 Experimental setup and conditions
Figure 2: Sketch of the experimental setup, with measurements of bubbles, drops, and dry aerosol particles. Bubbles of nearly identical sizes are generated by compressed air flowing through needles at the bottom of the tank. When two underwater pumps are turned on (as illustrated in the sketch), the injected bubbles are broken up into many smaller bubbles as they rise into the region of turbulence generated by the water jets from the pumps (bubbles not drawn to scale). The bubbles are imaged from the side view to measure the rising bulk bubbles, and from the top view to capture the bubbles at the surface. Large fields of view (black) and small fields of view (orange) are drawn approximately to scale and permit measurement of bubble radii from 30 µm to 5 mm. Liquid drops ejected into the air by bursting bubbles are measured through an inline holographic setup (measuring liquid drop radii from 10µm to 500 µm), while dry aerosol particles are measured by extracting and drying the liquid drops and analyzing them using an Optical Particle Sizer and Scanning Mobility Particle Sizer (OPS, SMPS composed of DMA and CPC - TSI, Inc.) (for dry particle radii from 20 nm to 5 µm). A line of filtered, compressed air is connected to the lid to flush the air above the water surface before the start of each run. Relative humidity and both air and water temperature are monitored throughout the experiments.
@@
The experimental setup shown in figure 2, consists of a polycarbonate tank with inner dimensions 47 x 47 x 61 cm3, similar to the one used by Néel & Deike (2021), which is filled to a height of 35 cm with a solution of artificial seawater (made of deionized (DI) water and artificial sea salt, ASTM D1141-98, Lake Products Company LLC). The water height is chosen so that the waves/fluctations on the water surface remain small and the surface bubbles can be clearly imaged. Bubbles are continuously generated as compressed air flows through the needles positioned at the bottom of the tank, arranged in a square with eight needles extending horizontally from each side. The air flowrate through the needles is controlled at 100 cm3/min per needle, so that the resulting plume of rising bubbles is fairly monodisperse.
@@
The artificial seawater solution is prepared by mixing DI water with ASTM D1141-98 Artificial Sea Salt, which contains the same proportions of elements found in natural seawater in amounts greater than 0.0004% by weight (Lake Products Company LLC). Solutions discussed in this paper were prepared with a salinity of 36 g/L (35 g/kg) to mimic the composition of ocean water. The solution has liquid viscosity $\mu$ = 1.001 mPa s, liquid density $\rho$ = 1025 kg/m3, and static surface tension $\gamma$ = 52 mN/m. The low static surface tension indicates that the salt mixture likely includes some surfactant. At the beginning and end of each day of experiments, the solution was characterized by measuring the surface tension isotherm with a Langmuir Trough (KSV NIMA, model KN 1003). This measurement was used to confirm that the surface properties of the solution remained consistent over multiple days of experiments. Two pumps (Rule 20DA 800 GPH Bilge Pump) are mounted 6 cm above the needle tips, pointed towards each other from opposite corners. When the pumps are turned on, they produce water jets that meet in the center of the tank, creating a region of turbulent flow that causes the rising bubbles to fragment into many smaller bubbles, with turbulence similar to the setup described in Ruth et al. (2022). Experiments can be performed with or without forced turbulence, where the turbulence has two effects: agitate the flow and break the bubbles, leading to a broad-banded size distribution. Note that bubble imaging occurs above the region of active breakup.
@@
Case Description Bubble Total Air Needle ID Pumps # Needles $R_{b,inj}$ $N_{R_{b}<2mm}$
Production Flowrate (µm) (mm) (cm-3)
Method (cm3/min)
Narrow - 20mm Needles 3200 432 off 32 2 0.40
Broad - 20mm Needles 3200 432 on 32 2 45
Broad - 20mm - half volume Needles 1600 432 on 16 2 25
Narrow - 30mm Needles 3200 965 off 16 3 0.13
Broad - 30mm Needles 3200 965 on 32 3 43
Narrow - 2500µm Porous Media 800 off 0.25 57
Table 1: Experimental parameters. The bulk bubble size distribution across cases was varied by changing between quiescent and turbulent flow underwater (by turning the underwater pumps off or on, causing the bubbles to break up in the turbulent case), by changing the injection bubble size (using two different needle sizes or an aquarium bubbler), and by halving the volume (using 16 instead of 32 needles with the same air flowrate per needle). Case names describe the range of radii over which a significant number of bubbles are concentrated (narrow, broad) and the initial injection bubble size before any breakup in turbulence (20mm, 30mm, 2500µm). The Broad-20mm-half volume case is characterized by bubbles arriving at the surface isolated from each other, bursting individually instead of clustered as the smaller number of needles leads to both reduced volume and increased needle spacing. $R_{b,inj}$ represents the typical injection bubble size by the needles or bubbler. $N_{R_{b}<2mm}$ shows the number of bubbles (per unit volume) smaller than radius 20mm, which reinforces the designation of certain cases as narrow- or broad-banded. We provide the total air flowrate through the needles for each case, but we caution that this flowrate value should not be used to approximate the bubble population in our configuration, as it is primarily related to the volume contained in the largest bubbles and does not capture the variations in the bubble size distributions (broad and narrow configurations) necessary when linking bursting bubbles to emitted drops. Bubble size distributions for each case are shown throughout sections 2.2.4, 4.1, and 4.2. Representative surface bubble images for each case are shown in figure 1.
@@
In order to capture the large range of length scales present in the system – spanning bubble radii of 30 to 5000 µm and drop sizes of 0.05 to 500 µm – we make multiple measurements of bulk bubbles, surface bubbles, drops, and dry particles. Bulk bubbles are imaged in both a large and small field of view by cameras positioned above the region of active breakup. Surface bubbles are also captured in two fields-of-view by cameras positioned perpendicular to the water surface, where the bubbles cluster near the center of the tank before bursting and ejecting drops into the air above the water surface, which is enclosed in the tank with an acrylic lid.
Drops are sized using a combination of an inline holographic system positioned just above the water surface (Erinin et al., 2019, 2023), an Optical Particle Sizer (OPS: TSI, Inc. OPS 3330), and a Scanning Mobility Particle Sizer (SMPS, consists of: TSI, Inc. Advanced Aerosol Neutralizer, Model 3088; Electrostatic Classifier, Model 3082; Differential Mobility Analyzer (DMA), Model 3081; Condensation Particle Counter (CPC), Model 3752). The SMPS operates with an aerosol flowrate of 0.3 L/min and sheath flowrate of 2.4 L/min. The OPS and SMPS are connected through the lid and pull air with suspended drops out of the tank, where they are dried (TSI Diffusion Dryer, Model 3062) and then measured. Relative humidity and air temperature are monitored 25 cm above the calm water surface throughout all experimental runs, along with the temperature of the solution.